# DynEx: Dynamic Code Synthesis with Structured Design Exploration for Accelerated Exploratory Programming

### v1

@@ 3.1. Design Exploration through the Design Matrix
Figure 2. DynEx’s Design Matrix User Interface: Users traverse through the matrix, which guides them through exploring the Person, Approach, and Interaction dimensions relevant to the problem space. For each dimension, users begin on the Idea level by inputting their problem (A). They then brainstorm ideas (B), and can iterate on the ideas (C). They can select an idea (C) and submit it in the input box (E). They then move on to the Grounding level, brainstorm a response (F), and submit that as well. They have the option to save versions (G), in case they want to move back to the Idea level and explore a new idea. Previous entries used to curate the current entry response are highlighted in yellow. Finally, users can can explore the prototype (H).
@@
A user begins by suggesting a problem they would like to solve, such as “I need reminders water my plants” or “I need to plan my meals”– it can be vague, something they have not thought of a solution for. Once submitted, the user is led through the Design Matrix, a structured framework that guides users through the design space. The matrix is centered on three key dimensions that are foundational to user-centric design, which make up the columns:
@@
    Interaction - the core user interaction paradigm, such as a table, card-swipe, news feed, or chat-bot layout.
Each dimension is explored on two levels of specificity, which make up the rows of the matrix:
@@
    Idea: the brainstorming of the higher-level concepts across each dimension
@@
    Grounding: the solidification of foundational details through iterative deepening that specifies how the idea should be developed. For example, if the user selected a card-swipe UI as the interaction paradigm, the grounding should specify what information exists on each card, what swiping right means, and what swiping left means.
Fig. 2 illustrates the Design Matrix, with columns for Person, Approach, and Interaction, and rows for Idea and Grounding. The Design Matrix facilitates the transformation of abstract concepts into actionable design components, allowing the user to fully understand their problem and solution space before settling on a prototype design.
3.1.1. Design Matrix Interaction
Users can traverse through the Design Matrix in any order, as long as they submit the Idea before the Grounding level for each dimension. On the Idea level of specificity, the system help users brainstorm different ideas through divergent thinking; the Grounding level of specificity solidifies the idea through convergent thinking. To reference entries in the matrix, we use a column:row notation; for example – Person:Idea, Approach:Grounding.
In order to ground the idea, the idea must be already submitted for that particular dimension; Person:Idea must be submitted before Person:Grounding, Approach:Idea must be submitted before Approach:Grounding, and Interaction:Idea must be submitted before Interaction:Grounding. Otherwise, the user can brainstorm the Person, Approach, and Interaction dimensions in any order. For example, if the user has the application idea of “tinder for groceries”, the interaction paradigm is clear – they want a card-swipe UI. They can fill out the Interaction dimension first before formalizing the Person and the Approach. If the user wants to build a journalling application to help with mental health that uses cognitive behavioral therapy (CBT) principles, they can fill out the Approach dimension to use CBT before filling out the Person and Interaction dimensions.
We use context curation to guide the outputs of each entry in the matrix. The system takes in all previously-submitted entries of the matrix, since they are already-known aspects of the project, to help inform the response of the current entry. The previous entries used as context are highlighted in yellow on the UI for the user to visualize (see Fig. 2 - I). If the user chooses to resubmit a dimension on the Idea level, the Grounding for that dimension will not be factored in the context. Responses for each entry were generated using few-shot examples and can be found in Appendix 8.
The next sections dive into the specifics of each matrix entry, using a motivating example to illustrate the process. The user inputs a problem they want to solve: learn Chinese.
@@ 3.1.2. Person
The user collaborates with the system to brainstorm who the application is for in Person:Idea. The system brainstorms potential target users and returns 3 results:
  1. (1)
    ”Non-native speakers interested in Chinese culture”
  2. (2)
    ”Visual learners struggling with language memorization”
  3. (3)
    ”Travel enthusiasts planning a trip to China”
Each of these cases have nuances; for (1) the user would probably be more interested in creating an app to learn Chinese culture rather than Chinese words; for (2), the user prefers a visual learning application; for (3), the user would probably focus more on learning travel-related vocabulary and phrases.
The user can generate as many ideas as they want, iterate on them and re-brainstorm, or directly update the input box. For example, if the user wanted more results that focused on learning Chinese culture, they could type that in the iteration box and brainstorm ideas along that direction. The user chooses to brainstorm more ideas, and 3 more ideas are added to the list:
  1. (4)
    ”Retired adult wanting to expand linguistic skills”
  2. (5)
    ”Busy university student wanting to study Chinese in his free time”
  3. (6)
    ”Chinese-born American looking to brush up language skills”
All of these results have different implications as well. For (4), the retired adult presumably has more time to learn Chinese – their method for learning would be different than (5), the busy university student who can only study at short intervals. Additionally, (6) presents a different use-case, a user who already has existing Chinese knowledge. Ultimately, the user selects (2) ”Visual learners struggling to learn with language memorization”. The user moves on to Person:Grounding to iteratively deepen their idea.
When grounding the Person dimension, the system aims to fully understand the user goals, challenges, and the broader context in which the problem should operate. Person:Grounding also defines the specific problems that users face, some shortcomings of existing solutions, and how the application would address these gaps. It takes in Person:Idea as context, which turns yellow on the interface. The system presents a list of bullet points that deepen the idea:
  • ”Confusion arises due to unfamiliarity with Chinese characters and their complex structure, slowing the learning process”
  • ”Difficulty in linking characters to their corresponding meaning or pronunciation, hindering vocabulary acquisition”
  • ”Traditional memorization methods offer little aid to visual learners who could better recall information through imagery”
The user can directly edit, save this version, or iterate on it. Once they are satisfied, they can submit their result and move on to the Approach dimension.
@@ 3.1.3. Approach
In the Design Matrix entry for Approach:Idea, the system brainstorms methods or strategies that the application could employ to solve the problem, given the existing context of the matrix. Since Person:Idea and Person:Grounding were submitted, those entries are highlighted yellow and used to inform the Approach:Idea response. The system presents 3 Approach ideas:
  1. (1)
    ”Pictorial spaced repetition learning”
  2. (2)
    ”Visual storytelling for language acquisition”
  3. (3)
    ”Cognitive load theory for efficient memorization”
The user can generate as many ideas as they want, iterate on them and re-brainstorm, or directly update the input box. The user is intrigued by (1) ”Pictorial spaced repetition learning”, and (2) ”Visual storytelling for language acquisition”. They select (2) ”Visual storytelling for language acquisition” and move on to Approach:Grounding to explore more details.
In Approach:Grounding, the system concentrates on the practical implementations by defining the essential components and logic required of the chosen approach. It takes in the existing matrix as context - Person:Idea, Person:Grounding, and Approach:Idea. For (2) ”Visual storytelling for language acquisition”, the system returns:
  • ”Integrate visual aids, such as illustrations or animations, that correlate with each word’s meaning to enhance understanding and recall”
  • ”Develop an algorithm that links related words and images together in a meaningful story, promoting stronger memory associations”
  • ”Make use of GPT to generate context-rich sentences or mini-stories, helping to create a narrative around each word or character”
  • ”Ensure design of the learning material caters to visual learners, with a focus on vibrant, engaging, and contextually relevant graphical representations”
The user feels inspired by this, and saves it using the “save version” button in case they want to come back to it (see Fig. 2 - G). The user goes back to Approach:Idea and selects (1) ”Pictorial spaced repetition”, then moves back to Approach:Grounding. The system generates this:
  • ”Implement the spaced repetition system (SRS) algorithm to schedule review times according to each user’s progress, allowing items to reappear before they’re likely to be forgotten”
  • ”Incorporate visual images representing Chinese characters or words, enhancing the memory association between visual cues and corresponding meanings”
  • ”Integrate regular reviews into the learning process, which are fine-tuned according to the user’s performance, to further solidify memory and retention”
  • ”Leverage cognitive science principles such as interleaving (mixing similar tasks) and retrieval practice (recalling an item from memory), to increase learning effectiveness”
The user decides (1) ”Pictorial spaced repetition” is more suitable for their current needs, and moves on to the Interaction dimension.
@@ 3.1.4. Interaction
In the Design Matrix entry for Interaction:Idea, the system brainstorms the high-level interaction paradigm of the application, such as a swipe interface or chatbot. In this example it takes in the Person and Approach dimensions as context. The system returns
  • ”Simple guess-and-review quiz interface”
  • ”Visual dictionary flashcard interface”
  • ”Image-based language learning game interface”
The user selects (1) ”Simple guess-and-review quiz interface” as the idea and moves on to Interaction:Grounding. In Interaction:Grounding, the system specifies the features and details of the UI components. It determines what information will be displayed, the nature of user interactions, and how these elements will contribute to the overall user experience. The grounding is provided:
  • ”The quiz interface should present the Chinese character or word along with its corresponding image”
  • ”The user then attempts to guess its meaning. If they respond accurately, the item is pushed back into the review cycle based on the SRS algorithm”
  • ”If the guess is incorrect, the correct meaning is displayed, and the item is scheduled for another review sooner”
  • ”Users should have a clear view of their progress and a way to navigate to previously learned words for self-study”
The user is happy with this and submits it. Now that the Design Matrix has been traversed, they can begin implementation in code.
3.2. Dynamic Iterative Implementation With LLM Code-Generation
@@
DynEx leverages LLM code-generation capabilities to create the prototype. DynEx self-invokes multi-modal LLMs to enable the creation of a rich suite of applications. While prototypes should be simple, they need a minimum level of complexity and detail in order to properly emulate the user experience. Applications with images should be prototyped with images, applications that require data need relevant data, and applications that require dynamic data, such as personalization or recommendation apps, should be able to replicate that. Because our generated code calls generative AI, we enable the creation of rich applications that can sufficiently prototype the user experience.
The system first identifies the project requirements, creates the project specification (spec), and generates synthetic placeholder data if required (see Fig  3- C, D E). Once the initial setup is complete, the user and system collaborate in creating the prototype through step-by-step implementation, where the system breaks down the implementation plan into steps (see Fig  3- F). Because project specifications are often complex and detailed, attempting to complete the application with a one-shot approach often results in key elements being missed, leading to non-functional prototypes. With missing functionalities, users can spend unnecessary time debugging and adding features to align with the initial spec. Breaking down the spec into manageable steps ensures that details are not missing in the final prototype. Furthermore, our iterative step-by-step approach saves the code at each stage, providing a natural form of version control.
3.2.1. Project Requirements
Before implementation, the system identifies prerequisite project requirements and generates a (spec) (see Fig  3- C, D, E). The project requirements detail whether or not the prototype requires GPT, images, placeholder data, or supported external libraries. The system brainstorms these requirements using the Design Matrix as context. Users can manually update and change the project requirements if needed - for example, if the application is a book recommendation app, and DynEx did not recommend the images as a requirement, the user can add that requirement themselves if they so please. Using the project requirements and Design Matrix as context, the system then generates a project specification to translate the abstract ideas across each dimension into a technical implementation.
Self-invoked Multi-Modal LLMs
Our project requirements include the ability to call GPT and generate images. To do this, we introduce a technique to self-invoke multi-modal LLMs, enabling the creation of a diverse suite of applications. This allows our code to generate dynamic data, create images, and even return code through LLMs. In software development, it is common to rely on third-party application programming interfaces (APIs) to enhance functionality. For example, travel websites can call the Google Flights API for flight information, and restaurant recommendation applications can call the Yelp API for restaurant data. Generative AI APIs are among the most powerful and versatile, excelling at handling a wide variety of tasks. By self-invoking multi-modal LLMs, we can build applications that are otherwise not supported by static data; for example, book recommendation applications that require dynamic data from LLMs, wardrobe visualization applications that can call image models, and note-organizational applications that can use LLMs to categorize notes and sensemake. LLMs are a powerful tool to self-reference when generating code for exploratory programming. It allows us not only streamline the development process and rapidly prototype applications, but also create more dynamic and functional UIs.
Dynex self-invokes GPT-4 and DALL-E 2 to enable dynamic and flexible content creation based on the project requirements. We use few-shot examples to support this technique; if GPT or Images is selected as a requirement (see Fig  3- C), we provide different code examples of calling GPT-4 or DALL-E 2 in the back-end prompt to guide the self-invocation. If these requirements are not needed, we omit those examples. For our example to learn Chinese, ”images” is selected – DynEx will self-invoke DALL-E 2 to generate images for the prototype.

### v3

@@ 3.1. Design Exploration through the Design Matrix
Figure 2. The Design Matrix: There are three columns for the Person, Approach, and Interaction dimensions. There are two rows for the levels of specificity: Idea and Grounding. The matrix can be traversed in any order as long as the Idea comes before the Grounding for each dimension. Each entry will take all already-submitted entries as context to generate a response. For example, the Interaction:Grounding will take in all previous entries as context in this diagram.
@@
A diagram of DynEx’s Design Matrix. There are three columns for the Person, Approach, and Interaction dimensions. There are two rows for the levels of specificity: Idea and Grounding. The matrix can be traversed in any order as long as the Idea comes before the Grounding for each dimension. Each entry will take all already-submitted entries as context to generate a response. For example, the Interaction:Grounding will take in all previous entries as context in this diagram.
The Design Matrix is a structured framework that facilitates the transformation of abstract concepts into actionable design components. It is a crucial part of exploratory programming, allowing users to explore their problem and solution space and balance trade-offs between design choices before settling on a prototype design. The Design Matrix is rooted in the dimensional approach to design space exploration. The matrix organizes complex design challenges into three distinct dimensions to ensure that all these facets are thoroughly considered and addressed before application development. Three key design dimensions make up the columns:
@@
    Interaction - the interaction paradigm that the user interface is implemented around. For example, a card-swipe, news feed, or table layout.
Each dimension is explored on two levels of specificity, which make up the rows:
@@
    Idea - a phrase that conveys the main concept of a dimension. For example, in the Interaction dimension, one idea could be ”flashcards implemented in a card-swipe UI.”
@@
    Grounding - 3-5 bullet points that give concrete details for how the idea is realized. For example, the grounding for the card-swipe idea should specify what information is displayed on each card, what action is associated with swiping right, and what action is associated with swiping left.
The Person dimension describes who the user is, what needs they have, and why they are creating this application. In human-centric design, understanding a person’s context and motivations is key to creating an application truly tailored to user needs. For example, potential users of a language learning application might range from a business traveler to a student interested in cultural immersion. A business traveler prioritizes practical, situational language skills tailored to business communication. In contrast, a student interested in cultural immersion prioritizes conversational fluency and deeper engagement with aspects such as history, pop culture, and cuisine.
The Approach dimension describes the underlying mechanism to achieve the user’s goal; it details the theory, algorithm, or model that is the essence of the solution to the problem. For example, in a learning application, learning theories that can drive an application include spaced repetition, a learning technique to review information at gradually increasing intervals to enhance long-term memory retention, and gamification, a technique that uses game elements to encourage learning.
The Interaction dimension describes how information should be displayed and engaged with in the UI. For example, for a flashcard application, a card-swipe interaction complements a spaced repetition approach because it provides a natural, dynamic way for users to move through cards at their own pace. Other potential interactions that pair with spaced repetition could include a calendar-based interface to schedule learning reminders, or even a news-feed interface with quizzes.
In the Design Matrix, each dimension has two levels of fidelity: Idea and Grounding. The idea is a low-fidelity concept that is simple to understand, but lacks precision. The grounding is a higher level of fidelity that expands the idea, providing details that specify how to fully realize this idea in a prototype. It paints a more comprehensive vision of the design.
For each entry in the matrix, LLMs generate suggestions the user can accept, edit, or regenerate. The Design Matrix uses all its previously-submitted entries to make suggestions for the current entry. This happens both across rows, where the grounding depends on the idea, and across columns. This ensures that the current design direction is considered when brainstorming suggestions for each entry and that ideation across dimensions is both consistent and relevant, ultimately leading to a cohesive design solution. For example, if the Approach (ie: spaced repetition) is already filled out, the system will suggest a corresponding Interaction (ie: flashcard swipe). Together, these 6 entries describe the prototype design in enough detail that it can be implemented in code.
3.2. Code-Generation through Modular Stepwise Implementation
3.2.1. Project Scoping
Before implementing the prototype, DynEx must (1) identify technical requirements and (2) generate a project specification (spec). DynEx uses the contents of the Design Matrix to identify technical requirements. If the prototype needs to support data visualization and flow charts, external libraries, such as ChartJS and GoJS, can be used. If the prototype needs data generated on-the-fly, such as for personal recommendations, the prototype will call GPT. If images are required, the prototype will call DALL-E 2. Identifying technical requirements prior to development minimizes ambiguity and missteps during implementation.
DynEx uses the technical requirements and the Design Matrix to generate a detailed, bullet-point spec, typically about one page long. Specifically, the spec details the application layout (what content will preside in the interface), the user interaction (how the application will respond when a user clicks on buttons), and logic of the application (when and how the application will call LLMs or use external libraries). It outlines the exact functionalities of the application.
3.2.2. Modular Stepwise Implementation
Figure 3. DynEx Code Generation: The system first decomposes the project into steps that are modifiable by the user. The user iteratively implements each step. The user can add, remove, or update steps at any point. The backend LLM then generates the UI code – the user can also interact with the UI . The user can directly modify the code output, debug, and revert to previous versions. The user iterates through steps until the final prototype is complete.
[…]
A diagram of DynEx’s code generation process. The system first decomposes the project into steps that modifiable by the user. The user iteratively implements each step. The user can add, remove, or update steps at any point. The back-end LLM then generates the UI code – the user can also interact with the UI . The user can directly modify the code output, debug, and revert to previous versions. The user iterates through steps until the final prototype is complete
To implement the prototype, DynEx employs a modularized, stepwise approach to development. It breaks down the spec into a multi-step plan, where each step is the next-smallest testable unit of the previous step. At each step, users test and approve the code before moving on to the next until the completion of the prototype. They can also edit, remove, and add steps at any point in the implementation process.
DynEx generates a multi-step plan because specs are complex and detailed, and attempting to generate code with a one-shot approach often results in key elements being missed. Each step is modular and nonoverlapping, similar to how developers incrementally adds features that build on prior progress. The first step establishes the basic HTML structure. The following steps add functionality in a logical order, allowing the system to generate code iteratively without needing to rework or disrupt the code before it. For example, when creating a book recommendation application, the process starts by setting up the HTML structure, building different sections of the application, and finally tool-calling GPT.
The modular steps also provide version control, a necessary functionality in exploratory programming systems. Since each step is saved, DynEx provides a clear version history that captures the most advanced working prototype while offering the flexibility to iterate between versions. Users can always fall back on the last successful version, ensuring that they always have a stable point to return to.
3.2.3. Self-invoked Multi-Modal LLMs
Tool-calling is a known method to make LLMs more powerful by having them interact with external APIs. In DynEx, the generated code can call the OpenAI API to generate data and images as needed. For example, a book recommendation system needs to call an LLM after a user has selected preferences to get recommendations. An outfit recommender system will call image models to generate previews of the user’s outfits based on their existing wardrobe. We call this method self-invoking multi-modal LLMs. It enables the creation of powerful and realistic applications that can generate data to respond to user intent, a capability not offered by prototypes supported by static data.
3.2.4. Creating Multiple Prototypes
Once a prototype is created, users can create additional variations by revisiting the Design Matrix, tweaking parameters to explore variations, and comparing the results. Users can modify entries in the Design Matrix and re-implement new designs to generate a diverse set of prototypes. This is a consistent method with exploratory programming, empowering users to explore multiple solutions in parallel and refine their designs.
4. Usage Scenario
To illustrate the user experience of DynEx, we present the example of a user, Franky. Franky wants to create a personalized application to learn Chinese that addresses his specific learning needs, but does not have a clear design in mind. He starts by typing ”learn Chinese” in the ”Problem” box, and clicks ”Submit” (Fig 4 - A).
4.1. Design Exploration through the Design Matrix
Figure 4. DynEx’s Design Matrix User Interface: Users traverse through the matrix, which guides them through exploring the Person, Approach, and Interaction dimensions relevant to the problem space. For each dimension, users begin on the Idea level by inputting their problem (A). They then brainstorm ideas (B), and can iterate on the ideas (C). They can select an idea ( D) and submit it in the input box (E). They then move on to the Grounding level, brainstorm a response (F), and submit that as well. They have the option to save versions (G), in case they want to move back to the Idea level and explore a new idea. Previous entries used to curate the current entry response are highlighted in yellow (I). Finally, users can can explore the prototype (H).
[…]
The Design Matrix appears on the screen. He sees three columns for the Person, Approach, and Interaction dimensions, and two rows for Idea and Grounding – the two levels of specificity. To reference entries in the matrix, we use a column:row notation; for example – Person:Idea, Approach:Grounding.
@@ 4.1.1. Person
To choose an the idea for the Person (Person:Idea), Franky clicks the ”Brainstorm” button(Fig 4 - B). After a moment, three potential target users appear in that entry (Fig 4 - D).
1. Non-native speakers interested in Chinese culture 2. Visual learners struggling with language memorization 3. Travel enthusiasts planning a trip to China
Franky can generate as many ideas as he wants (Fig 4 - B) or iterate on them and re-brainstorm (Fig 4 - C). Curious to see more suggestions in the Person dimension, Franky clicks the ”Brainstorm” button again, and 3 more ideas are added:
4. Retired adult wanting to expand linguistic skills 5. Busy university student wanting to study Chinese in his free time 6. Chinese-born American looking to brush up language skills
Franky selects (2) on the interface – he identifies as a visual learner, recognizing that many of his learning challenges stem from lack of visual aids. ”Visual learners struggling with language memorization” shows up in the input box (Fig 4 - E), which he can directly update if he wants. He accepts the idea as is, excited to see more details about this in the grounding row (Person:Grounding). A list of bullet points that motivate why a visual learner might want this application appear in the grounding entry on the interface:
Confusion arises due to unfamiliarity with Chinese characters and their complex structure, slowing the learning process. Difficulty in linking characters to their corresponding meaning or pronunciation, hindering vocabulary acquisition. Traditional memorization methods offer little aid to visual learners who could better recall information through imagery.
This reasoning resonates with Franky, especially the second bulletpoint. As a visual learner, his struggle with traditional text-based language learning stems from the lack of meaningful visual connections that aid memory. Franky has the option to directly edit the grounding, save this version, or iterate on it, but he accepts the grounding as is, excited to build a learning application with a visual approach. He moves on to explore the Approach column in the matrix.
@@ 4.1.2. Approach
Franky generates ideas for the Approach dimension (Approach:Idea) :
1. Pictorial spaced repetition learning 2. Visual storytelling for language acquisition 3. Cognitive load theory for efficient memorization
Franky is interested in exploring both ideas (1) and (2). He selects (2) first and moves on the ground this idea. In the grounding entry (Approach:Grounding), DynEx details an approach that connects the Chinese phrases together to generate a narrative and correlates each phrase to an image within the story. Franky appreciates these details – he hadn’t considered visual storytelling as an approach before. He then selects (1) ”Pictorial spaced repetition learning” and grounds the idea. The grounding provides concrete details to the spaced repetition algorithm, such as pairing each image to a word and scheduling review times according to the user’s progress. After seeing the groundings for both ideas (1) and (2) (the full outputs can be seen in Appendix I), Franky decides to implement (1) ”Pictorial spaced repetition learning”, since he believes he is better suited to learning the Chinese phrases through memorization, rather than storytelling. He moves on to the Interaction dimension.
@@ 4.1.3. Interaction
Franky generates ideas in the Interaction dimension (Interaction:Idea). Three ideas appear on in the matrix entry: quiz, flashcard, and game interface, all interaction paradigms that effectively showcase a spaced repetition algorithm. Franky selects the quiz idea – he doesn’t want a game interface, thinks flashcards aren’t engaging enough, but believes that quizzes are a quick and effective approach to test language acquisition. He moves on to ground the idea in the final entry of the matrix (Interaction:Grounding). The grounding specifies how the quiz interface will be formatted. The quiz will present the Chinese phrase along with its corresponding image and prompt the user to guess its English meaning. The prototype will allow users to view their progress and self-study before the quiz. The full output for the Interaction column can be seen in Appendix I. Franky thinks that the grounding covers all the essential features he needs in a learning application: a way to study, test, and track progress. He names the prototype (Fig 4 - H) and submits the matrix. Now that the Design Matrix has been traversed, he can begin implementation in code.
4.2. Code-Generation through Modular Stepwise Implementation
@@
4.2.1. Project Scoping
To scope the project out before implementation, Franky clicks the ”Identify Project Requirements” button (Fig 5 - C). After a moment, the checkboxes for ”Dynamically generated AI-images” and ”Pre-generated data” are selected. This makes sense to him – the pictorial approach he specified in the Design Matrix requires images, and data needs to be generated to populate the Chinese phrases he will learn. Franky has the option to check or uncheck requirements if he wants, but he is satisfied with what DynEx has recommended. Franky then clicks ”Generate New Spec” (Fig 5 - D), and a spec appears, detailing that the data must contain Chinese words, phrases, and IDs, and specifying that images must be returned for each quiz question to assist with visual learning. It describes 3 sections: Flashcards for self learning, Quiz for knowledge testing, and Progress for the user to track their learning journey (Appendix J). Finally, Franky clicks ”Generate Data” (Fig 5 - E), and Chinese phrases to practice with appear on the screen (Appendix K). Franky can regenerate or directly edit the outputs for the spec and the data, but he is satisfied with the overall project scoping. He moves on to generate code for the prototype.
4.2.2. Modular Stepwise Implementation
Figure 6. Step Outputs: UI outputs created at each step in the usage scenario.
[…]
Screenshot of UI outputs created at each step of the usage scenario.
To begin implementation, Franky clicks ”Generate Plan” on the interface, and a step list breaking the spec into multiple steps pops up on the screen (Fig 5 - F):
  1. (1)
    Set up the React application and create the main layout with the three sections: ’Flashcards’, ’Quiz’, and ’Progress’
  2. (2)
    Implement the ’Flashcards’ section
  3. (3)
    Implement the ’Quiz’ section, utilizing the spaced repetition algorithm
  4. (4)
    Implement the ’Progress’ section
  5. (5)
    Call GPT to generate images for the Chinese characters
Franky can regenerate, approve, or modify this step list, but he likes this plan and starts to generate code for each step sequentially. He clicks step (1) ”Create the HTML structure” to navigate to that part of the implementation phase and clicks the ”Generate Code” button (Fig 5 - H). After a moment, the prototype’s UI is rendered for Franky to interact with. He sees a basic HTML, and in the Flashcards section, there is a grid filled with the pre-generated Chinese phrases (Fig 6 - Step 1). He clicks the ”Quiz” and ”Progress” tab – the content for these tabs are empty, but since this step only created the HTML structure, that is expected. He can also view or edit the code by clicking the ”Code” tab (Fig 5 - K). Franky can either accept this code or fix it – he accepts it and proceeds to the next step.
Franky clicks (2) ”Implement the ‘Flashcards section’” and generates the code for that step. In the Flashcards section, he can now click left and right to swipe through the cards and study the Chinese phrases (Fig 6 - Step 2). He approves step 2, and moves on to step 3.
After generating code for (3) ”Implement the ‘Quiz’ section,” Franky discovers that there’s an error with the “submit” button in the Quiz section. The Quiz section shows the Chinese phrase and prompts the user to type in the corresponding English meaning as expected, but after the user submits their answer, the prototype displays both the next quiz question and the next quiz question’s answer. Franky can debug this by iteratively prompting the system in the right direction to regenerate code (Fig  5 - I), redoing the step (Fig  5 - H), or clicking the ”Code” tab (Fig 5 - K) to debug the generated code by hand. Franky makes use of the “Iterate” box and types in the input: “After I click submit, it moves on to the next question but also shows the answer of the next question.” After re-generating the code, the debugged UI and code is shown on the interface. Franky can iterate as many times as he wants and switch between versions until he is happy with the final step output (Fig  5 - J). Franky tests the regenerated code – the Quiz section is working as expected, so he approves step 3 and moves on to step 4.
Franky generates the code for (4) ”Implement the ‘Progress’ section”. In the prototype’s UI, he sees a metric for accuracy rate in the Progress section. When he tests the interface, he sees his accuracy rate go up based on how he scores on the Quiz section. Pleased that he can now track his progress, he moves on to step 5.
After generating the code for (5) ”Call GPT to generate images.”, Franky sees that the Quiz section now has images alongside each Chinese phrase. Now, he has visuals to assist him in language learning. The prototype for the original implementation plan is complete.
Franky tests the prototype and decides that he would like another feature: in the Progress section, Franky would like a visual indicator to show which words he has mastered. He clicks the ”Add Step Beneath” button (Fig  5 - G), and types in ”Allow users to see Mastered words” in the input box. He then generates the code for this new step and tests the prototype – all the phrases he has answered correctly in the Quiz section now show up in the Mastered Words section under the Progress tab.
After mastering the initial pre-generated phrases in the prototype, Franky wants a way to expand his initial dataset. He adds a step, types ”suggest new phrases for me to learn,” and generates the code for that step. Now, after Franky masters his initial set of words, the prototype will use tool-calling to dynamically suggest new phrases with GPT, ensuring that he always will have novel and relevant phrases to learn whenever he wants. Satisfied that he can continuously expand his language skills, Franky decides his prototype is complete.
5. System Implementation
The system was implemented in Python, Typescript, and Flask. The Design Matrix uses GPT-4 — due to its vast understanding across various domains, it is well suited for brainstorming and refining ideas. DynEx use Claude 3.5 Sonnet for code generation — Claude is known to be better at writing UI code, and it enables us to use Claude Artifact as a baseline for comparison in our evaluation. Claude is also highly responsive to explicit instructions, which is crucial for code-generation. The repository is open-sourced on Github. [^1] https://github.com/jennygzma/ui-design-prototype
5.1. Design Matrix
When the user submits a problem, they use the system to brainstorm each entry in the Design Matrix. For each of the six entry types, we provide few-shot examples to guide the type of response desired. Three different applications are used as few-shot examples: OKCupid, Tinder, and Coffee Meets Bagel. These applications broadly span the range of the design space for dating apps; while they are all meant for single people, they have different target users, core approaches, and interaction paradigms. Each application example is broken down across dimensions and levels of specificity to establish the desired level of detail for each entry. The full few-shot examples can be seen in Appendix A.
Additionally, the system takes in all previously-submitted entries of the matrix, since they are already-known aspects of the project, to help inform the response of the current entry. The previous entries used as context are highlighted in yellow on the UI for the user to visualize. If the user chooses to resubmit a dimension on the Idea level, the Grounding for that dimension will not be factored as input context. Ultimately, this ensures that the current design is always factored in when suggestions for each entry, and that the dimensions are synthesized to form a coherent prototype design.
5.2. Code Generation
5.2.1. Project Scoping
The system identifies the technical requirements using the Design Matrix as context, where the contents of the Design Matrix are converted into a natural language prompt detailing what each entry in the Matrix is. The technical requirements and Design Matrix are then used as context to generate the spec. We use few-shot examples to guide the generation, which can be found in Appendix B. We use the spec, along with few-shot examples, to generate the data, which can be seen in Appendix C.